# Designing Adaptive User Interfaces for mHealth applications targeting chronic disease: A User-Centric Approach

### v1

@@ 3. Methodology
We developed an AUI prototype of the mHealth application and gathered input from various users with chronic disease through a combination of focus group sessions, interviews, and surveys. Through the analysis of the data collected, we derived guidelines for effective AUI design for mHealth app targeting chronic disease. Subsequently, we refined these guidelines based on the feedback of the participants, as shown in Figure 2.
3.1. Adaptive User Interface prototype
Building upon insights from an earlier Systematic Literature Review (SLR) (Wang et al., 2023b), our research enhances the adaptation categories and integrates them into a single prototype that includes three primary types of adaptation: presentation adaptation, content adaptation, and behaviour adaptation. Details of the prototype can be found in (Wang et al., 2023a).
Presentation adaptation is a crucial aspect of UI design, aiming to enhance user experience by modifying the UI appearance. This adaptation involves adjusting elements such as colour, positioning, and font size to better suit the user’s preferences and needs. It is divided into two main categories: 1) graphic design: manipulates visual aspects, like theme, and layout; 2) information architecture: pertains to the structural organisation of information within the system.
Content adaptation is the process of adapting interface content to better suit user needs and preferences, encompassing various elements such as text, semantic elements, images and explanatory text. We propose two primary categories: 1) content complexity: simplifies content for better comprehension, which comprises three subcategories. The easy-to-understand language employs clear and concise language to convey information in a straightforward manner (see Figure 3(a)). Minimalist design streamlines the presentation of content by minimising visual clutter and extraneous details. Text-to-image conversion converts complex textual information into easily understandable visual representations, such as diagrams or illustrations, to aid comprehension, particularly for users who may struggle with textual content (See Figure 3(b)). 2) interface elements rearrangement: includes two main subcategories: position adjustments and selective element concealment. Position adjustments optimise the placement of interface elements to ensure intuitive navigation for users. The concealment of selective elements strategically hides or reveals interface elements based on user context or preferences, thus simplifying the interface and focusing on key content areas (see Figure 3(c)).
Behaviour adaptation involves adjusting navigation methods, enabling/disabling interface elements, and modifying interaction modes within an application. This complex form of adaptation often spans multiple steps and can encompass content and presentation adjustments. We group it into five subcategories: 1) navigation adaptation: alters user navigation permissions or suppressing specific modules; 2) add-on functions: incorporates new features to enhance user assistance and application usability; 3) pervasive strategies, adapts motivation techniques for effective behavioural changes tailored to user types/status; 4) multimodal interaction: adjusts interface modalities based on varied usage contexts; and 5) difficulty level adjustment: modifies game/exercise difficulty to balance engagement and challenge based on user motivation/performance. The AUI prototype was evaluated by three coauthors and tested on two real users with experience with mHealth apps. For each adaptation, we provided transitions, sample rationales, and additional instructions to facilitate user navigation and interaction with various adaptations.
3.2. User study design
Our user study consists of two parallel investigations, as illustrated in Figure 2. We conducted a qualitative investigation via focus groups and interviews to examine how individuals experience AUIs in the context of chronic diseases using the AUI prototype. At the same time, we conducted a quantitative survey to collect user preferences regarding different dimensions of adaptations.
@@ 3.2.1. Focus group and interview studies
We designed a focus group and interview protocol consisting of three sections. The first section collected detailed demographic information about participants, including their chronic disease and their usage pattern of mHealth apps through a Qualtrics survey. The second section collected the participant’s views on the different adaptations we present in the AUI prototype. The participants initially reviewed a brief adaptation video accompanied by audio explanations, providing detailed introductions to each adaptation. This visual aid was followed by hands-on interaction with the prototype. Instructions were provided in the accompanying slides, encouraging participants to actively engage with the prototype. In the event of any difficulties, participants could refer to the instructions for assistance, ensuring a seamless user experience. The researcher remained readily available to offer support, allowing participants to focus solely on the example adaptations. Given the interpersonal dynamics inherent in the interview and focus group settings, participants may tend to offer responses that they perceive as socially desirable (Harris and Brown, 2019). This potential bias is further compounded by the qualitative nature of the data analysis process, which is based on inductive coding and is susceptible to the subjectivity of the researcher (Hoda, 2021). We adopted between-method triangulation to address inherent flaws in each method. This involved the triangulation of data using a combination of quantitative and qualitative techniques (Fusch et al., 2018; Denzin, 2017; Oppenheim, 2000). In the third section, participants in the interview and focus group study were required to complete a post-session survey to express their preferences for various adaptations explored with the AUI prototype during the session. This post-session survey consists of 7 questions. Further details of the post-session survey are provided in Appendix A. As a token of appreciation, all participants were offered a AU$30 virtual gift voucher.
Both the focus group and the interview sessions were subjected to one round of pilot test that involved varying numbers of participants. We employed a theoretical sampling method (Hoda, 2021) for participant recruitment. This sampling approach involves selecting new participants based on specific criteria as data collection and analysis progress and yield categories and concepts, allowing the deepening or expansion of existing categories and concepts. Qualitative data collection and analysis followed an iterative process spanned three different data collection iterations (see Table 1). Upon obtaining ethics approval from the university ethics board, 111Monash Human Subject Ethics Committee approval number: 36568 we started our focus group study by reaching out to potential participants through a mailing list associated with a previous chronic disease research programme. In this iteration, our main focus was on recruiting older individuals (Iteration 1). Subsequently, using our personal connections and networks, we expanded our participant pool to include several young adults with chronic diseases (Iteration 2).222The first author, who is afflicted by a chronic condition, engaged with both online and offline support groups for individuals in similar circumstances. To broaden the inclusion of participants, we also used the snowball method (Parker et al., 2019) during Iteration 2. To improve the reliability and validity of the data through multiple data collection methods (methodological triangulation) (Denzin, 2017), we employed semi-structured interviews alongside focus group studies in Iteration 3. The group dynamics inherent in focus group research provides a unique perspective for observing collective attitudes and shared perceptions (Acocella, 2012). At the same time, interviews allow for personalised exploration of individual experiences and perspectives, fostering a detailed understanding of nuanced perspectives (Wilson, 2014). Interview participants were recruited through a multifaceted approach. Using local community channels, recruitment efforts were expanded to include prominent healthcare facilities such as the Baker Institute Diabetes Clinic and Alfred Hospital (Iteration 3).333https://kidney.org.au, https://strokefoundation.org.au, https://www.dementia.org.au, https://www.alfredhealth.org.au, https://baker.edu.au/health-hub/clinics/diabetes-clinics
3.2.2. Survey
Our post-session survey aims to gather additional data about user perspectives regarding our prototype and to triangulate them with the findings from the focus groups and interviews. In parallel, we conducted an anonymous online survey using Google Forms aimed at individuals with chronic diseases to gather feedback on their preferences regarding various aspects of collecting user’s preference for different aspects of adaptation. These two surveys differ in their question designs and serve for different purposes. The survey consisted of four main sections, each focused on specific topics to ensure participants’ ease of understanding: demographic information, health status, mHealth application use patterns, and AUI preferences. To ensure the participants’ understanding of AUIs, a brief explanation with two examples of AUIs in the health domain was provided at the beginning of the AUIs section. This section focuses on participants’ preferences regarding various adaptations, different types of data, data collection methods, and the desired level of participation in the adaptation process. The survey questions were informed by a previous SLR (Wang et al., 2023b), and the survey went through expert reviews and pilot tests with five participants. The survey data collection occurred from January to July 2023 through various recruitment channels, with a total of 90 respondents (Figure 2). The survey data collection went through three phases: 1) email ads and social media (58 responses), 2) physical posters distributed at Baker Institute Diabetes Clinic and Alfred Hospital (13 responses), and 3) advertisements through Dementia Australia, Stroke Foundation, and Kidney Australia (19 responses). Participants who completed the survey were eligible for a AU$20 virtual gift card draw. You can find the survey instruments in the Appendix B.
3.3. Guidelines Development
Our guideline development process follows and adapts the framework proposed by Hermawati and Lawson (2015) and Quiñones et al. (2018). We use data collected during survey, interview, and focus group sessions, as well as the existing literature on guidelines, design considerations, and criteria, to establish the guidelines. By synthesising this qualitative and quantitative information, we were able to identify common categories, patterns, and user needs that served as the basis for our guideline development. In addition, our guideline specification involved a process of refinement and validation, where we refined the guidelines based on the feedback of the participants. After the refinement process, we generated nine guidelines that serve as a practical and effective resource for designing adaptive mHealth applications targeting chronic diseases.
@@ 3.4. Data analysis
3.4.1. Qualitative data analysis
@@
We used the data analysis procedures of Socio-Technical Grounded Theory (STGT) (Hoda, 2021) to analyse the data from the focus group and interview study. This decision was primarily driven by two key factors. First, the focus of our study closely aligns with the principles of the socio-technical research framework that STGT is built upon. Our investigation revolves around AUIs in applications related to chronic diseases, a socio-technical phenomenon that encompasses both human and social aspects (such as chronic disease, ageing, and computer literacy), as well as technical aspects. Second, STGT allows for selective application by integrating its core data analysis procedures of open coding (with hashtags) and constant comparison, along with memoing, within a mixed-method research study. Traditional grounded theory methods such as Glaserian (Glaser, 1992) and Strauss-Corbinian (Strauss and Corbin, 1994) are developed as standalone methodologies for theory development. By leveraging STGT for data analysis, we adopted a holistic approach to data analysis, recognising the complex and interconnected nature of the socio-technical phenomena under investigation. We obtained consent from the participants to transcribe the audio recordings, and subsequently stored and analysed the data using NVivo.
@@
The data collection and analysis process followed an iterative and interleaved approach, as depicted in Figure 4. Saturation, which indicates the point at which no new categories or concepts properties emerged, was achieved by Iteration 3. The qualitative data was analysed by the first author and was subsequently shared with the remaining authors to encourage collaborative discussion at each stage of the process. This collaborative effort helped refine the analysis and determine the optimal presentation of the findings in all iterations of data collection. In Section 4, we present the key concepts and categories derived from the STGT analysis. An example of process of applying STGT for data analysis (Hoda, 2021) is provided below:
@@
Open Coding and Constant Comparison: We analysed the audio transcripts and extracted various codes from the raw data. We provide some examples below.
@@
Raw Quote: ”If you want to set it, then you don’t want to change it. Likewise, you don’t want to have to go through all those settings again. I think that would be complicated, and I am not sure I could handle it on my own. ” (FP 4)
@@
Code 1: mental workload for adaptationRaw Quote: “I am not going to log my daily blood sugar levels in the app because I am too lazy to do it sometimes. Will the adaptation hide this important function because I don’t use it a lot?” (FP 6)
Code 2: user preferences contradict the app’s intended usageConcept: The two code examples given above suggest a concept: ”The user may not be the right person to handle the adaptation”
@@
Category: Drawing insights from the memos generated during the coding process and the codes and concepts identified, we derived various categories. The given example illustrates the category concerning who should take charge of the adaptation process.
@@
Theoretical sampling: We started our initial convenience sampling with the primarily elderly in three focus group studies Iteration 1. After we conceptualised concepts from the focus group data, we applied theoretical sampling to obtain more data that helped us to explicate the emerging concepts. For example, most of the focus group participants, who are older, express difficulty engaging in the adaptation process and prefer it to occur without their direct involvement, as the multistep nature of the process can be confusing for them. Conversely, a participant of a much younger age has a contrasting view, showing interest in adapting the features or functions. As a result, the category ”The user may not be the right person to handle the adaptation” . needs to be refined, prompting us to recruit more participants from a younger age group in Iteration 2 (see Figure 4).
@@
Memoing played a crucial role in our approach, allowing us to explore emerging concepts and potential relationships between them, as outlined by Hoda (2021). These memos served as invaluable tools for capturing key insights and reflections gleaned from our open-coding efforts, which are further detailed in Section 4.5.

### v2

@@ 3. Methodology
We conducted the study in two distinct stages to systematically develop and validate guidelines for AUI design in mHealth applications that target chronic disease management. In stage one, we developed the AUI prototype tailored for chronic disease-related applications and conducted qualitative research through interviews and focus group studies with individuals managing chronic diseases. In parallel, a quantitative survey was administered to capture user preferences related to different aspects of adaptation. In stage two, the analysis of both qualitative and quantitative findings, combined with insights synthesized from existing literature, led to the development of an initial set of guidelines for designing adaptive mHealth applications. The preliminary guidelines are further evaluated and refined through survey feedback from both end-users and software practitioners, resulting in the finalization of nine guidelines. The refined guidelines are then validated by applying them to real-world mHealth applications. We also compared our guidelines with existing mHealth usability guidelines and analyzed whether the issues identified through our approach aligned with those mentioned in user app review (see Figure 2).
3.1. Stage One: User Study Design
Building on the insights from an earlier SLR (Wang et al., 2023b), our research extends the adaptation categories and integrates them into a prototype that includes three primary types of adaptations: presentation adaptation, content adaptation, and behavior adaptation. Details of the prototype can be found in Wang et al. (2023a).
@@ 3.1.1. Focus group and interview studies
Grounded in the SLR on contemporary developments in AUI within the field of chronic diseases (Wang et al., 2023b), we designed a focus group and interview protocol consisting of two sections. The first section collected detailed demographic information about the participants, including their chronic disease and their use pattern of mHealth applications through a Qualtrics survey 111https://www.qualtrics.com. The second section collected the participant’s views on the different adaptations we present in the AUI prototype. The participants initially reviewed a brief adaptation video accompanied by audio explanations, providing detailed introductions to each type of adaptation. This visual aid was followed by hands-on interaction with the prototype. Instructions were provided in the accompanying slides, encouraging participants to actively engage with the prototype. In the event of any difficulties, participants could refer to the instructions for assistance, ensuring a seamless user experience. The researcher remained readily available to offer support, allowing participants to focus solely on the example adaptations. As a token of appreciation, all participants were offered an AU$30 virtual gift voucher. Employing a theoretical sampling approach to recruit participants entails choosing new individuals based on particular criteria, which include the advancement of data collection and analysis, as well as the development of categories and concepts (Hoda, 2021). Qualitative data collection and analysis followed an iterative process in three different data collection iterations, with information on the recruiting iteration presented in Wang et al. (2024). To improve the reliability and validity of our findings through methodological triangulation (Denzin, 2017), we employed both semi-structured interviews and focus groups. Focus groups provided valuable insight into collective attitudes and shared perceptions through dynamic group interactions (Acocella, 2012), while interviews allowed a detailed exploration of individual experiences, allowing for a more nuanced understanding of personal perspectives (Wilson, 2014). By combining these methods, we captured both the breadth of group consensus and the depth of individual viewpoints.
3.1.2. User survey
Participants in interviews and focus groups may provide socially desirable responses, introducing bias (Harris and Brown, 2019), which is further compounded by the subjectivity of inductive coding in qualitative analysis Hoda (2021). To mitigate these biases, between-method triangulation was employed, involving the triangulation of data using a combination of quantitative and qualitative techniques (Fusch et al., 2018; Denzin, 2017; Oppenheim, 2000). We conducted an anonymous online survey using Google Forms aimed at individuals with chronic diseases to gather quantitative feedback on their preferences on various aspects of adaptation. The survey design is informed by the SLR conducted in AUI in the context of chronic diseases (Wang et al., 2023b). The survey data collection went through three phases: 1) email ads and social media (58 responses), 2) physical posters distributed at Baker Institute Diabetes Clinic and Alfred Hospital (13 responses), and 3) advertisements through Dementia Australia, Stroke Foundation and Kidney Australia (19 responses). Participants who completed the survey were eligible for an AU$20 virtual gift card draw. The user survey questions are provided in the Appendix A.
3.2. Stage Two: Guidelines Development and Evaluation
The guideline development process follows and adapts the framework proposed by Hermawati and Lawson (2015) and Quiñones et al. (2018). Data collected through surveys, interviews, and focus group sessions in stage one, along with findings from existing literature on guidelines, design considerations, and evaluation criteria, were synthesized to inform guideline creation. The process also included iterative refinement and validation based on participant feedback, ensuring the relevance and applicability of the guidelines.
3.2.1. End-user guideline evaluation survey
We carried out a preliminary evaluation of our guidelines with end-users, primarily concentrating on assessing their clarity and perceived usefulness. Considering the potential scarcity of documentation on design considerations and guidelines, especially in emerging technology domains, it is imperative to maintain user involvement throughout the process, and ensure that guidelines, though originally created for developers, designers, and researchers, hold significance for end-users (Hermawati and Lawson, 2015; Zaphiris et al., 2007; Lechner et al., 2013; Zaphiris et al., 2009). Therefore, we strived to validate the efficacy and acceptance of these guidelines with the same cohort of users involved in our interview and focus group study (Stage One). We reached out to all participants via email, inviting them to complete a guideline evaluation survey. Participants were briefed on the findings of the user study and provided with an overview of the generated guidelines. To ensure anonymity and simplify the process, participants used a unique withdrawal code from their initial registration to link their responses to the feedback survey and user study. Subsequently, participants were requested to evaluate each of the guidelines and offer suggestions regarding additions, removals, or edits for each guideline.
3.2.2. Software practitioners guideline evaluation survey
After end-user evaluation, another online survey was conducted among software practitioners and other relevant stakeholders involved in mHealth application development to evaluate the effectiveness of the proposed guidelines. In the survey, respondents were provided with definitions of mHealth applications and AUI, along with access to the proposed guidelines via publicly available links. Demographic information was gathered from respondents, along with their evaluations of the proposed guidelines. Participants were specifically asked to assess the applicability of the guidelines in real-world practice, and to provide insights into their strengths, limitations, and potential areas for improvement. The survey questions for evaluating the guidelines, which are adapted from Shamsujjoha et al. (2024), can be found in Appendix B. Following ethics approval, we conducted a pilot study with representatives from the target population to evaluate the survey’s design, specifically the clarity of the questions and response options. Overall, participants expressed satisfaction with the survey and did not suggest changes to the questionnaire structure or answer formats. However, they recommended the inclusion of comprehension checks to ensure that the respondents had adequately understood the guidelines, especially in cases where the guidelines may not be thoroughly reviewed prior to answering related questions. The survey was revised to include two comprehension check questions, allowing participants two attempts to complete them, with exclusion from the study if both attempts were failed. Moreover, participants who do not meet the screening criteria, which focused on their expertise in developing health-related applications, particularly those targeting chronic diseases, were also excluded. Traditional attention check questions were not included, as the study emphasized qualitative responses and prioritized meaningful engagement and understanding of the proposed guidelines. An initial set of 9 responses was collected through personal networks. Due to the limited sample size, survey distribution was subsequently expanded via Prolific222https://www.prolific.com/, resulting in 34 additional responses and a total of 43 participants. During recruitment on Prolific, a customized screening tool was employed to ensure that participants were professionals working in the technology sector. To maintain the integrity of the sample, individuals whose responses did not align with their prescreening information on Prolific were excluded from the study.
3.2.3. Case study
[…]
@@ 3.3. Data Analysis
3.3.1. Stage One: Qualitative data analysis
@@
We used the data analysis procedures of Socio-Technical Grounded Theory (STGT) (Hoda, 2021) to analyze data from the focus group and the interview study. This decision was primarily driven by the close alignment between the focus of our study and the principles of the socio-technical research framework that STGT is built upon, as our investigation revolves around AUIs in applications related to chronic diseases, a socio-technical phenomenon that encompasses both human and social aspects, as well as technical aspects. STGT allows for selective application by integrating its core data analysis procedures of open coding, constant comparison, and memoing, while traditional grounded theory methods such as Glaserian (Glaser, 1992) and Strauss-Corbinian (Strauss and Corbin, 1994) are developed as standalone methodologies for theory development. We obtained consent from the participants to transcribe the audio recordings, and subsequently stored and analyzed the data using NVivo333https://lumivero.com/products/nvivo/.
@@
The data collection and analysis process followed an iterative and interleaved approach (Figure 3). Saturation, which indicates the point at which no new categories or concepts properties emerged, was achieved in the third iteration of the user study. The qualitative data was analyzed by the first author and was subsequently shared with the remaining authors to encourage collaborative discussion at each stage of the process. In Section 4, we present the key concepts and categories derived from the STGT analysis. An example of a process for applying STGT for data analysis is provided below.
@@
Open Coding and Constant Comparison: We analyzed the audio transcripts and extracted various codes from the raw data.
@@
Raw Quote 1: ”If you want to set it, then you don’t want to change it. Likewise, you don’t want to have to go through all those settings again. I think that would be complicated. ”
Code 1: mental workload for adaptation
@@
Raw Quote 2: “I am not going to log my daily blood sugar levels in the application because I am too lazy to do it sometimes. Will the adaptation hide this important function because I don’t use it a lot?”
@@
Code 2: user preferences contradict the app’s intended usageThe two code examples given above suggest: Concept: The user may not be the right person to handle the adaptation. Drawing from the memos generated during the coding process (see Figure 4), along with the identified codes and concepts, it illustrates the Category: Who should take charge of the adaptation process.
@@
Theoretical Sampling: Initial data collection began with convenience sampling, primarily involving older adults in focus group sessions. After we conceptualized concepts from the focus group data, theoretical sampling was employed to further refine and elaborate on the emerging themes. For example, most of the focus group participants in Iteration 1, primarily older adults, express difficulty engaging in the adaptation process and prefer it to occur without their direct involvement, as the multi-step nature of the process can be confusing to them. In contrast, a participant of a much younger age has a contrasting view, showing interest in adapting the features or functions. As a result, the concept the user may not be the right person to handle the adaptation needs to be refined, prompting us to recruit more participants from a younger age group in Iteration 2 (see Figure 3).
@@
Memoing played a crucial role in our approach, allowing us to explore emerging concepts and potential relationships between them, as described by Hoda (2021). These memos served as invaluable tools for capturing key insights and reflections gleaned from our open-coding efforts, which are further detailed in Section 4.2. An example memo is provided in Figure 4.
3.3.2. Stage Two: Qualitative data analysis
Given the relatively small amount of unstructured data collected in stage two survey studies, which contrasts with the richer data from stage one focus group and interview studies, we do not anticipate discovering multi-layered findings (Hoda, 2024). Consequently, thematic analysis is employed in stage two to delve deeper into the qualitative aspects of the survey responses, allowing the discovery of common themes that surpassed the survey responses (Braun and Clarke, 2012, 2006). The qualitative data was analyzed by the first author and was subsequently shared with the remaining authors to encourage collaborative discussion.